OpenStudy (anonymous):

The sum of the first 7 terms of an arithmetic series is 329. The common difference is 14. a) Find the value of the first term. b) 69,800 is the sum of the first n terms of the sequence. Find n. The sixth term of an arithmetic sequence is 49 and the fifteenth term is 130. a) Find the common difference for this sequence. (I got this, it's d=9) b) Find the first term. c) How many f the terms of this sequence have a value which is less than 300?

OpenStudy (holly00d1248):

329/7 is your first term

OpenStudy (anonymous):

@Holly00d1248 Gotcha. So 47

OpenStudy (holly00d1248):

yep

rishavraj (rishavraj):

@dellzasaur nah....how it could be 47....huh

OpenStudy (anonymous):

@rishavraj He said to divide 329/7 to get the first term for the first problem

rishavraj (rishavraj):

nah its not the way.....u need to find the first term???i mean u need help in part a??

rishavraj (rishavraj):

u remember sum of "n" terms = (n/2)*[2a + (n-1)d] a - first term d - common difference

OpenStudy (anonymous):

@rishavraj Yeah

rishavraj (rishavraj):

plug the values and get "a" i.e first term....:)

OpenStudy (anonymous):

@rishavraj I got -13/2? I'm not sure

rishavraj (rishavraj):

don't take it otherwise but how u managed to get a negative value ...that also fraction...lol see sum of n terms = 329 n = 7 d= 14 now solve

rishavraj (rishavraj):

\[329 = (7/2)*[2a + (7 - 1)14]\]

OpenStudy (anonymous):

@rishavraj 35?

rishavraj (rishavraj):

nope a = 5.....solve it again..... its 94 = 2a + 84 get "a"....remember "a" is the first term

OpenStudy (anonymous):

@rishavraj Ok I got it now

rishavraj (rishavraj):

@dellzasaur need help in other parts too???

OpenStudy (anonymous):

@rishavraj Please! I really need help on b of the first question and b and c of the second question. I think I may be able to do b on the second one but i just want to make sure

rishavraj (rishavraj):

@dellzasaur see now its 69800 = (n/2)*[2a + (n - 1)d] a = 5 d = 14 get "n"...:))

OpenStudy (anonymous):

@rishavraj Thank you :) I'm working it out but I'm not sure what to do with the n's in the equation. Can you show me one or two steps of it so I can see? I can probably go from there!

OpenStudy (anonymous):

Would I divide everything by 2?

OpenStudy (anonymous):

or multiply the 69800 by 2? I dont think that's right

rishavraj (rishavraj):

u will get fking quadratic equation....:(( \[69800 = (n/2)*[10 + (n - 1)14]\]

rishavraj (rishavraj):

@dellzasaur see for question no 2 its \[a _{6} = a + (n - 1)d\]

rishavraj (rishavraj):

and \[a _{15} = a + (15 - 1)d\]

rishavraj (rishavraj):

\[a _{6} = a + (5 - 1)d\] u having two equation and two variables...get "a" and "d"

OpenStudy (anonymous):

@rishavraj the d is 9 right?

OpenStudy (anonymous):

I got 13 and 4 for a, not entirely sure

rishavraj (rishavraj):

@dellzasaur in question 2 common difference will be 9....and u first term = 4...as u said....